# Quantization-Aware Neuromorphic Architecture for Efficient Skin Disease Classification on Resource-Constrained Devices

### v1

@@ Abstract
Accurate and efficient skin lesion classification on edge devices is critical for accessible dermatological care but remains challenging due to computational, energy, and privacy constraints. We introduce QANA, a novel quantization-aware neuromorphic architecture for incremental skin lesion classification on resource-limited hardware. QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference. Its quantization-aware head and spike-compatible transformations enable seamless conversion to spiking neural networks (SNNs) and deployment on neuromorphic platforms. Evaluation on the large-scale HAM10000 benchmark and a real-world clinical dataset shows that QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% Top-1 accuracy and 81.7% macro F1 on the clinical dataset, consistently outperforming leading CNN-to-SNN models under fair comparison. Deployed on BrainChip Akida hardware, QANA achieves 1.5 ms inference latency and 1.7 mJ energy per image, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, and exceeding the performance of advanced CNN-to-SNN conversion methods. These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
Index Terms:
Resource-Constrained Devices, Edge Computing, Neuromorphic Computing, Low Latency, Energy Efficiency.
@@ I Introduction
In this work, we present an end-to-end pipeline for efficient and incremental skin lesion classification on neuromorphic hardware. As illustrated in Fig. 1, our approach consists of four main stages: (1) data preprocessing, including image quality screening, augmentation, and SMOTE-based class balancing; (2) construction of a quantization-aware neural network with stacked Ghost modules, channel attention, and a spike-compatible output head; (3) automated conversion of the trained network to a spiking neural network (SNN) using the Akida MetaTF toolkit; and (4) deployment and optimization on the BrainChip Akida neuromorphic platform for low-latency, energy-efficient inference. The core QANA architecture is specifically designed for hardware compatibility, enabling direct and lossless translation to spike-based event-driven computation. This pipeline directly addresses the major challenges of edge deployment, including computational constraints, class imbalance, limited data, and the need for incremental model updates.
Experimental results on both the HAM10000 public benchmark and a real-world clinical dataset demonstrate the practical advantages of our approach. QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% / 81.7% on the clinical dataset. On Akida hardware, the system delivers 1.5 ms inference latency and 1.7 mJ energy consumption per image, surpassing state-of-the-art CNN-to-SNN conversion baselines in both accuracy and efficiency. These findings confirm the effectiveness of QANA for resource-constrained, real-time medical image analysis.
@@ II Related Work
Prior studies on neuromorphic hardware and Spiking Neural Networks (SNNs) have highlighted their suitability for edge inference and continual learning in constrained environments [14]. Platforms such as BrainChip’s Akida [15], IBM’s TrueNorth [16], and Intel’s Loihi [17] demonstrate ultra-low-power operation and support on-chip incremental learning, enabling model updates without full retraining. These features are advantageous for medical imaging tasks on edge devices. However, existing works primarily focus on general benchmarks like MNIST [24] or CIFAR [25] and rarely address domain-specific challenges in dermatology [26]. Furthermore, in medical image contexts, edge Spiking Neural learning methods often suffer from catastrophic forgetting and fail to preserve performance on previously learned classes when new samples are introduced [27].
CNN-to-SNN conversion has become a common strategy to leverage mature CNN architectures—ResNet, DenseNet—for deployment on neuromorphic processors [28]. Conversion toolkits, including Akida’s CNN2SNN [29], deliver compatibility by quantizing weights and replacing activation functions with spike-based equivalents. While some studies report preserved accuracy on standard datasets [30], performance significantly degrades on small or imbalanced medical datasets due to batching overfitting, quantization noise, and inability to capture rare lesion features [15]. Limitations in mapping CNN operations—such as batch normalization, global average pooling, and multi-bit activations—to spiking neurons further compound accuracy loss during conversion [31].
In parallel, conventional CNN-based methods for skin lesion classification, such as those using DenseNet-121 [32] or Inception-v4 [33], show high performance on benchmark datasets like ISIC. Despite their effectiveness, these models necessitate continuous retraining on cloud infrastructure to address domain shifts, new lesion types, or demographic variations [18]. Such retraining incurs latency, high computational cost, and data privacy concerns [34]. Moreover, general CNN models often fail to generalize to underrepresented conditions [35], such as rare tumors or images from diverse skin tones, highlighting the need for adaptive and privacy-preserving on-device solutions. These limitations motivate the development of neuromorphic frameworks that maintain diagnostic performance while supporting incremental learning and efficient operation in resource-constrained settings.
@@ III-A1 Image Preprocessing
All dermatoscopic images are first screened for quality. Corrupted, low-resolution, or artifact-contaminated images are excluded using automated checks based on image metadata and pixel statistics, followed by manual review when ambiguity remains. For compatibility with the Akida neuromorphic platform, all accepted images are resized to $64\times 64$ pixels using bilinear interpolation, which is the maximum supported resolution under current on-chip memory constraints. Pixel intensities are normalized channel-wise to the range $[0,1]$ to ensure consistent input distribution. No denoising or inpainting operations are performed, as quality control steps eliminate samples with significant artifacts.
@@ III-A2 Data Augmentation
To mitigate overfitting and increase the diversity of the training dataset, each image in the training split undergoes stochastic augmentation with the following operations: brightness adjustment (random factor sampled uniformly from $[0.7,1.3]$), contrast modification ($[0.8,1.2]$), random horizontal and vertical flipping (each with probability 0.5), hue shift (random shift in $[-0.08,0.08]$ in HSV space), and saturation variation ($[0.85,1.15]$). Each augmentation is applied independently to each sample with the specified probability. All augmentation parameters are determined by fixed random seeds for reproducibility. These transformations simulate real-world imaging variation while preserving lesion semantics.
@@ III-A3 Synthetic Minority Oversampling (SMOTE)
To address class imbalance, especially in rare lesion categories, the Synthetic Minority Oversampling Technique (SMOTE) [36] is applied to the training data. Given a minority class dataset $\mathcal{S}=\{x_{1},x_{2},...,x_{n}\}$ in $\mathbb{R}^{d}$, for each $x_{i}\in\mathcal{S}$, the $k$ nearest neighbors $\mathcal{N}_{i}=\{x_{i_{1}},...,x_{i_{k}}\}$ are identified in feature space via Euclidean distance. For each synthetic sample, a random neighbor $x_{i_{j}}$ is selected, and a new sample is generated by:
$x_{\text{new}}=x_{i}+\lambda\cdot(x_{i_{j}}-x_{i}),\quad\lambda\sim\mathcal{U}(0,1)$ (1)
where $\lambda$ is independently sampled for each feature dimension. Repeating this process for all $x_{i}$ and multiple neighbors results in the augmented set:
$\mathcal{S}_{\text{aug}}=\mathcal{S}\cup\left\{x_{i}+\lambda^{(m)}\cdot(x_{i_{j_{m}}}-x_{i})\mid x_{i}\in\mathcal{S},\,1\leq m\leq M\right\}$ (2)
where $M$ is the number of synthetic samples generated per original sample. All images are maintained at $64\times 64$ resolution to conform with Akida’s memory requirements throughout oversampling. SMOTE is applied only to the training set to preserve test set integrity and eliminate information leakage.
@@ III-B2 Spike-Compatible Feature Transformation
$\widehat{F}=\min\left(1,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (7)
@@ III-C CNN-to-SNN Conversion for Neuromorphic Deployment
The conversion of a trained convolutional neural network (CNN) into a spiking neural network (SNN) is a critical step for enabling event-driven inference on neuromorphic hardware. In this work, we utilize the Akida MetaTF toolkit [41, 15] to perform an automated and quantization-aware transformation of the CNN backbone described above into an SNN model suitable for direct deployment on the BrainChip Akida processor.
III-C1 Conversion Principles and Workflow
The Akida conversion [15] process follows a structured pipeline to ensure hardware compatibility and the preservation of model accuracy:
  • Operator Mapping: Each supported CNN layer (e.g., convolution, batch normalization, separable convolution, etc) is mapped to its spiking equivalent. For instance, ReLU activations are replaced by thresholding mechanisms that convert analog outputs to binary spike events. Layers incompatible with SNN operation (such as global average pooling or flatten) are substituted with spike-compatible alternatives, such as local pooling or spike-generating readout heads.
  • Quantization: All network weights and activations are quantized to a limited bit-width (8-bit), matching the precision constraints of the neuromorphic hardware. The quantization parameters are derived during the training and conversion phases to minimize information loss.
  • Temporal Spike Encoding: Continuous-valued activations from the CNN are converted into spike trains through rate coding or threshold-based event generation. The spike generation logic ensures that temporal information is preserved and the event-driven computation paradigm of SNNs is fully leveraged.
  • Resource and Constraint Adaptation: The MetaTF converter analyzes the input model to partition layers and neurons across the available hardware resources (neural cores, memory blocks) of the Akida chip, optimizing for parallelism, latency, and energy efficiency.
III-C2 Integration with Custom Backbone
Our proposed Quantization-Aware Network Architecture is specifically designed for seamless conversion. All intermediate activations are explicitly bounded and quantized, with network modules implemented in a form directly supported by Akida’s conversion pipeline. This ensures that no critical feature transformation is lost and that the functional mapping from input images to class predictions remains consistent between the CNN and its SNN counterpart.
III-C3 Conversion Output and Verification
Upon completion, the converter produces a deployable SNN model in Akida format. Model equivalence is empirically verified by comparing the output distributions of the original CNN and the converted SNN on a validation set. Any observed accuracy drop is mitigated through fine-tuning or incremental retraining on the SNN hardware, exploiting Akida’s support for on-chip learning.

### v2

@@ Abstract
On-device skin lesion edge classification remains hindered by the lack of efficient architectures compatible with neuromorphic deployment, as conventional CNN-to-SNN conversion suffers from structural incompatibility, performance degradation on rare diseases, and absence of continual learning. To address these limitations, we propose an end-to-end pipeline that incorporates a quantization-aware CNN backbone explicitly designed for conversion-preserving mapping to SNNs and neuromorphic deployment, culminating in the QANA architecture optimized for low-latency, energy-efficient, and incremental inference. QANA effectively integrates ghost modules, efficient channel attention, and squeeze-and-excitation blocks for robust feature representation with low-latency and energy-efficient inference. Its quantization-aware head and spike-compatible transformations enable seamless conversion to spiking neural networks (SNNs) and deployment on neuromorphic platforms. Evaluation on the large-scale HAM10000 benchmark and a real-world clinical dataset shows that QANA achieves 91.6% Top-1 accuracy and 82.4% macro F1 on HAM10000, and 90.8% Top-1 accuracy and 81.7% macro F1 on the clinical dataset, consistently outperforming leading CNN-to-SNN models under fair comparison. Deployed on BrainChip Akida hardware, QANA achieves 1.5 ms inference latency and 1.7 mJ energy per image, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, and exceeding the performance of advanced CNN-to-SNN conversion methods. These results demonstrate the effectiveness of QANA for accurate, real-time, and privacy-sensitive medical analysis in edge environments.
@@ I Introduction
We propose a quantization-aware neuromorphic architecture (QANA) within an end-to-end pipeline (Fig. 1) to overcome these challenge. QANA comprises three essential modules: (1) iterative feature extraction and downsampling block utilizing stacked Ghost modules, spatially-aware efficient channel attention (SA-ECA), and residual connections to reduce computation and inference latency, thereby directly facilitating effective CNN-to-SNN conversion; (2) spike-compatible feature transformation block employing quantization-aware normalization, bounded activations, and squeeze-and-excitation (SE) blocks, ensuring hardware compatibility and efficient incremental learning; and (3) quantized output projection block explicitly optimized for conversion-preserving CNN-to-SNN mapping and low-power neuromorphic deployment. These components collectively address key conversion limitations, preserve high accuracy for rare diseases, and efficiently support incremental model updates on edge devices.
Evaluations on the HAM10000 benchmark dataset show that QANA achieves a Top-1 accuracy of 91.6% and a macro F1 score of 82.4%. Further validation on a clinical dataset confirms robustness, achieving a Top-1 accuracy of 90.8% and a macro F1 score of 81.7%. When deployed on the Akida neuromorphic hardware platform, QANA attains an inference latency of 1.5 ms per image and energy consumption of 1.7 mJ, reducing inference latency and energy use by over 94.6%/98.6% compared to GPU-based CNNs, surpassing state-of-the-art CNN-to-SNN conversion methods in computational efficiency. These outcomes verify QANA’s practical suitability for resource-constrained and privacy-sensitive medical diagnostic applications.
Our main contributions are:
  • Conversion-preserving architecture. We design QANA with bounded activations, separable and ghost convolutions, and lightweight channel attention with SE gating, ensuring all operators lower one-to-one to 8-bit spiking kernels.
  • Deployment-aligned training and conversion. We provide a graph-preserving flow with batch-norm fusion and affine 8-bit quantization, combined with QAT, embedding-space SMOTE with safety filtering, and a lightweight on-chip incremental readout update, all under a fixed $64{\times}64$ input and matched preprocessing.
  • Edge evaluation. On Akida and under identical preprocessing and quantization for baselines, QANA achieves 91.6% Top-1 and 82.4% macro F1 on HAM10000 and 90.8%/81.7% on a clinical set, with 1.5 ms latency and 1.7 mJ per image.
@@ II Related Work
This section surveys the landscape for edge dermatology. We first outline neuromorphic hardware capabilities and constraints that govern feasible model design. We then compare ANN-to-SNN conversion with direct SNN training, including spike encoding and sparsity control, and finally review CNN-based dermatology systems as accuracy baselines. Across these threads we emphasize limitations under small and imbalanced datasets and strict latency and energy budgets, which together define the problem setting.
II-A Neuromorphic Hardware for Edge Imaging
Neuromorphic processors execute event-driven computation with low power and support on-chip adaptation, which matches the requirements of on-device medical imaging [14]. Akida, TrueNorth, and Loihi provide hardware support for spiking inference and incremental learning [15, 16, 17]. At the same time, these platforms constrain model design through limited operator sets, fixed-point quantization, and resource budgets that restrict input resolution and intermediate state. Related methods must therefore account for quantization and operator compatibility during both training and deployment.
II-B ANN-to-SNN Conversion
A prevalent strategy trains a conventional CNN and converts it to an SNN by mapping activations to spikes with calibrated thresholds and quantized ranges [24]. Conversion enables reuse of mature backbones but introduces several sources of error. Batch normalization and global pooling do not admit direct spike-based equivalents, which forces structural edits and alters feature statistics. Rate-coded readouts require temporal integration that increases latency and makes predictions sensitive to threshold and timestep calibration, which is difficult to tune on small and imbalanced datasets. Empirical studies report measurable accuracy drops after conversion under these conditions, indicating that quantization and operator mismatch are primary failure modes on edge workloads [15].
II-C Directly Trained SNNs and Spike Encoding
Direct training replaces the non-differentiable spike with surrogate derivatives and optimizes spiking dynamics end-to-end. SLAYER derives a time-domain error assignment rule for convolutional SNNs with learnable delays [25], and surrogate-gradient frameworks formalize stable optimization for deep networks [26]. Subsequent work improves few-timestep accuracy with residual spiking blocks [27] and temporal efficient training that couples the loss across discrete integration steps [28]. Despite these advances, training remains sensitive to surrogate choice, thresholds, and timestep schedules. Performance depends on spike encoding. Time-to-first-spike and related temporal codes reduce spikes and latency [29, 30], yet they complicate calibration on hardware designed for rate accumulation and may reduce robustness on fine-grained classes when data are scarce. Activity regularization controls event rates and energy [31, 25], but aggressive sparsity can underfit rare patterns. These factors make it difficult to guarantee accuracy, latency, and energy simultaneously under fixed hardware constraints.
II-D CNN-Based Dermatology Models
Dermatology classifiers based on DenseNet-121 and Inception-v4 reach strong results on ISIC-style datasets [32, 33]. In practice, deployment often depends on cloud retraining to handle distribution shift and rare categories, which introduces latency and privacy risk [18, 34]. Training data are class-imbalanced and images vary across devices and skin tones, which limits generalization to underrepresented conditions [35]. These limitations indicate that comparative evaluation on edge hardware should consider both conversion-based pipelines and directly trained SNNs with temporal encodings and sparsity control, so that the impact of quantization, operator compatibility, and encoding choices on accuracy, latency, and energy can be isolated under identical deployment settings.
@@ III-A1 Image Preprocessing
All dermatoscopic images undergo automated integrity checks followed by a brief manual review when needed. Inputs are processed at $64\times 64$ to match the Akida on-chip memory budget used in our deployment. To retain lesion content at this resolution, we apply Grad-CAM on the pretrained backbone to obtain a coarse saliency map. A tight square crop is centered on the top-activated region and resized with antialiased area resampling to $64\times 64$. If the saliency peak lies within 10% of the image center or the saliency map is flat (coefficient of variation <0.05<0.05), we fall back to a centered square crop. Pixel intensities are normalized per channel to $[0,1]$. No denoising or inpainting is used.
@@ III-A2 Data Augmentation
Augmentations preserve dermoscopic color cues. We use random horizontal and vertical flips (probability 0.5). Luminance is perturbed on the CIELAB $L$ channel with a multiplicative factor in $[0.95,1.05]$ or a gamma factor in $[0.97,1.03]$. Contrast is adjusted in $[0.95,1.05]$. No hue shift is applied. A gray-world normalization fitted on the training split is used during training only. All random draws use fixed seeds for reproducibility.
@@ III-A3 Synthetic Minority Oversampling (SMOTE)
We perform SMOTE strictly in the learned embedding space rather than on raw pixels. For each training image $x_{i}$ from class $c$, the backbone and spike-compatible transform produce a penultimate representation $\mathbf{r}_{i}\in\mathbb{R}^{4096}$ after global flattening. Let $\mathcal{I}_{c}$ be the index set of class $c$ and $\{\mathbf{r}_{i}\}_{i\in\mathcal{I}_{c}}$ the corresponding embeddings. Before neighbor search, embeddings are standardized on the training split only to avoid information leakage. A PCA whitening is then fitted on the same split and applied to stabilize distances in high dimension. Writing the sample covariance as $\mathbf{\Sigma}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ and retaining the top $d=256$ components, the whitened vector is
$\widetilde{\mathbf{r}}_{i}=\mathbf{\Lambda}_{d}^{-\frac{1}{2}}\mathbf{U}_{d}^{\top}(\mathbf{r}_{i}-\boldsymbol{\mu}),$ (1)
where $\boldsymbol{\mu}$ is the training mean, $\mathbf{U}_{d}\in\mathbb{R}^{4096\times d}$ the leading eigenvectors and $\mathbf{\Lambda}_{d}\in\mathbb{R}^{d\times d}$ their eigenvalues. Nearest neighbors are computed with Euclidean distance in the whitened space; for each $i\in\mathcal{I}_{c}$ we form a $k$-NN set $\mathcal{N}_{i}$ with $k=5$.
To reduce the influence of outliers, we keep only “safe” minority anchors whose local median neighbor distance is below the class-specific 90th percentile:
$\mathcal{S}_{c}^{\text{safe}}=\Big\{\,i\in\mathcal{I}_{c}\ \Big|$ $\operatorname{med}_{j\in\mathcal{N}_{i}}\big\lVert\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}$ (2)
$\leq Q_{0.9}\!\left(\big\{\operatorname{med}_{j\in\mathcal{N}_{u}}\big\lVert\widetilde{\mathbf{r}}_{u}-\widetilde{\mathbf{r}}_{j}\big\rVert_{2}\big\}_{u\in\mathcal{I}_{c}}\right)\Big\}.$
For each minority class $c$ with $N_{c}=|\mathcal{I}_{c}|$ and $N_{\max}=\max_{c^{\prime}}N_{c^{\prime}}$, we synthesize to a target level $\tau N_{\max}$ with $\tau=0.85$ to avoid distribution collapse. The number of synthetic embeddings is
$M_{c}=\max\!\big(0,\;\lfloor\tau N_{\max}-N_{c}\rfloor\big).$ (3)
Each synthetic vector is generated from an anchor $i\in\mathcal{S}_{c}^{\text{safe}}$ and a random neighbor $j\in\mathcal{N}_{i}$ by a convex interpolation and a small, locally aligned perturbation to improve diversity. Let $\mathbf{C}_{i}$ be the empirical covariance of $\{\widetilde{\mathbf{r}}_{u}\}_{u\in\{i\}\cup\mathcal{N}_{i}}$, and let $\mathbf{U}_{i}^{(m)}$ contain its $m=8$ principal directions. With $\delta_{i}=\operatorname{med}_{j\in\mathcal{N}_{i}}\|\widetilde{\mathbf{r}}_{i}-\widetilde{\mathbf{r}}_{j}\|_{2}$ and $\sigma_{c}=0.05\,\operatorname{med}_{u\in\mathcal{I}_{c}}\delta_{u}$, we sample
$\widetilde{\mathbf{r}}_{\text{new}}=\widetilde{\mathbf{r}}_{i}+\eta\big(\widetilde{\mathbf{r}}_{j}-\widetilde{\mathbf{r}}_{i}\big)+\sigma_{c}\,\mathbf{U}_{i}^{(m)}\boldsymbol{\xi},\quad\eta\sim\mathcal{U}(0,1),\;\boldsymbol{\xi}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{m}).$ (4)
The synthetic embedding is then mapped back to the original feature space by the inverse whitening:
$\mathbf{r}_{\text{new}}=\boldsymbol{\mu}+\mathbf{U}_{d}\mathbf{\Lambda}_{d}^{\frac{1}{2}}\widetilde{\mathbf{r}}_{\text{new}}.$ (5)
Synthetic data are not images. They are injected at the penultimate layer as additional feature vectors and used to train the quantized output projection with cross-entropy. During this stage the backbone up to the spike-compatible transform is frozen; only the classifier head is updated for 5 epochs with mixed real and synthetic embeddings. A subsequent short fine-tuning on real data only re-enables the last block and the head to reconcile the feature distribution. Input images are still resized to $64\times 64$ prior to feature extraction for hardware constraints, but no pixel-space mixing or image synthesis is performed at any time.
@@ III-B2 Spike-Compatible Feature Transformation
$\widehat{F}=\min\left(6,\max\left(0,\gamma_{\text{spk}}\cdot\text{BN}(\text{SepConv}_{3\times 3,256}(F^{(4)}))+\beta_{\text{spk}}\right)\right)$ (10)
@@ III-C CNN-to-SNN Conversion for Neuromorphic Deployment
We convert the trained quantization-aware CNN to an event-driven SNN using a graph-preserving flow with three stages: static reparameterization, integer quantization with calibration or QAT, and hardware-constrained code generation. First, all Batch Normalization layers are eliminated by fusing their affine statistics into the preceding convolution or pointwise projection. For a convolution output channel $c$ with kernel $\mathbf{W}_{c}$ and bias $b_{c}$ and BN parameters $(\gamma_{c},\beta_{c},\mu_{c},\sigma_{c}^{2},\epsilon)$, the fused weights and bias are
$\mathbf{W}_{c}^{\prime}=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\epsilon}}\mathbf{W}_{c},\qquad b_{c}^{\prime}=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\epsilon}}\left(b_{c}-\mu_{c}\right)+\beta_{c}.$ (13)
Depthwise and separable convolutions are treated identically channel-wise. Nonlinearities not in the target activation set are replaced during training by bounded ReLU6 so that all post-activation tensors are in $[0,6]$, which simplifies downstream spike generation and avoids undefined mappings in the spiking domain [40, 24]. Global average pooling is lowered to an integer spatial reduction with compile-time division embedded in the output scale. Flatten is treated as a metadata reshape and does not generate compute kernels. Softmax is applied on the host side after spike integration[24].
Quantization follows an affine, per-tensor, uniform scheme for both weights and activations with 8-bit fixed-point. Let $a$ denote a real activation and $(s_{a},z_{a})$ the scale and zero-point obtained by calibration on a held-out set or learned with quantization-aware training. The integer activation and its error are
$a_{q}=\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{a}{s_{a}}\right)+z_{a},\,0,\,255\right),e_{a}=a-\left(a_{q}-z_{a}\right)s_{a}.$ (14)
During SNN emission, each neuron integrates the quantized drive and generates a binary event whenever the internal accumulator crosses a programmable threshold $\tau$; resetting is subtractive. With state $u_{t}$ and drive $a_{q}$ the discrete-time dynamics are
$s_{t}=\mathbb{I}\!\left(u_{t-1}+a_{q}\geq\tau\right),\qquad u_{t}=\left(u_{t-1}+a_{q}\right)-\tau s_{t},$ (15)
and rate readout uses the spike count within an integration window. For ReLU6, $s_{a}=255/6$ and $z_{a}=0$ make the activation-range utilization explicit and fix the time-to-first-spike under a given $\tau$ [24].
Operator lowering is rule-based and deterministic. Convolution, depthwise convolution, separable convolution, average pooling, max pooling, elementwise clipping and integer linear layers are mapped one-to-one to supported spiking kernels. Global average pooling is lowered to an integer spatial reduction with compile-time division by the pooling area embedded in the output scale. Flatten is a metadata reshape that preserves memory layout and does not generate compute kernels. Unsupported nonlinearities such as Swish or hard-Swish are removed at training time and replaced by ReLU6. Squeeze-and-Excitation is realized by keeping the global pooling and two low-rank linear projections in the quantized graph and folding the final channel-wise gating into the subsequent pointwise projection when possible, otherwise emitting a per-channel integer rescale. Residual additions are integer tensor adds after scale alignment by rational multipliers that are absorbed into the successor scale. The converter rejects any subgraph that would require a floating-point dequantize. To make the comparison across backbones fair, a model is deemed convertible only if the compiled spiking graph contains no dequantize fallback and every tensor is quantized to 8-bit throughout the feature extraction and readout path; otherwise the model is excluded from neuromorphic evaluation and reported as a failed conversion.
Fairness metrics and logs are computed on the same train and test splits and under identical preprocessing. We retrain each CNN baseline with the same input resolution, the same augmentation policy, the same learning-rate schedule, and the same 8-bit quantization configuration, then convert and deploy the resulting SNN. Conversion success rate is defined on the quantized graph $\mathcal{G}$ as
$r_{\mathrm{succ}}=1-\frac{\lvert\mathcal{V}_{\mathrm{fp32}}\rvert}{\lvert\mathcal{V}\rvert}$, (16)
where $\mathcal{V}$ is the node set and $\mathcal{V}_{\mathrm{fp32}}$ are nodes that would require dequantization and thus cannot be placed on the neuromorphic cores. Accuracy drop is reported as $\Delta\mathrm{Acc}=\mathrm{Acc}_{\mathrm{SNN}}-\mathrm{Acc}_{\mathrm{CNN}}$ and $\Delta\mathrm{F1}=\mathrm{F1}_{\mathrm{SNN}}-\mathrm{F1}_{\mathrm{CNN}}$ on the same test set. For transparency, we release per-model conversion manifests listing, for every layer, the integer scale and zero-point, the applied reparameterization, and the final kernel selected by the code generator. The mapping rules in (13)–(16) and the practice of bounded activations to guarantee spike-compatibility follow established CNN-to-SNN conversion principles and 8-bit QAT literature.